OpenStudy (anonymous):

Studying for exam:

OpenStudy (anonymous):

\[(\sqrt{12}+\sqrt{-27}) - (\sqrt{3}-\sqrt{-48})\]

OpenStudy (anonymous):

Evaluate... I forgot how todo this.. it was in the start of the year

OpenStudy (anonymous):

\[\sqrt{12} = \sqrt{4 * 3} = 2 \sqrt{3}\] \[\sqrt{-27} = i \sqrt{27} = \sqrt{9*3} = 3\sqrt{3}\] \[\sqrt{-48} = i \sqrt{48} =i \sqrt{3* 16} = 4i sqrt{3}\]

OpenStudy (anonymous):

@boblovesmath I dont understand can you explain?

OpenStudy (cggurumanjunath):